What's the answer to x/6+2=4/5-x
Answer by algebrahouse.com(1659)

(x/6) + 2 = (4/5) - x
30[(x/6) + 2 = (4/5) - x] {multiplied entire equation by 30}
30(x/6) + 30(2) = 30(4/5) - 30x {multiplied each term by 30}
5x + 60 = 24 - 30x {cancelled and multiplied what's left}
35x + 60 = 24 {added 30x to both sides}
35x = -36 {subtracted 60 from both sides}
x = -36/35 {divided both sides by 35}

For more help from me, visit: www.algebrahouse.com